OpenStudy (anonymous):

determine whether y varies directly with x. If so, find the constant of variation k. 7y=-x

OpenStudy (saifoo.khan):

INverse.

OpenStudy (anonymous):

i'm gonna say k = 23

OpenStudy (anonymous):

or - 1/7 depending on what you meant,

OpenStudy (anonymous):

you would have to have solve for y on the left hand side \[y = -x/7\] So \[k = -1/7\]